Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
ruggi - 16:52 Tuesday 14 May 2024 (64254) Print this report
Recent unlocks to be understood

Besides the unlock due to an earthquake, and the ones intentionally or accidentally caused by someone, a small number of recent unlocks need to be explained. Here is the gps of the events:

1)  1398740055    2024-05-03 02:53:57

2)  1399357405    2024-05-10 06:23:07

3)  1399397637    2024-05-10 17:33:40

4)  1399453603    2024-05-11 09:06:25

5)  1399557370    2024-05-12 13:55:52

The only one for which the Lock Losses Monitor tries to give an explanation is #2: SSFS correction diverging

We see in fig 1 that SSFS Corr was actually increasing because a fast oscillation was growing below the main 150 Hz component. The value was still quite far from the saturation when the ITF unlocked: usually a standard SSFS oscillation can reach 10 V and then cause the unlock. Some disturbance was visible at the same time in some PSL and INJ signals: a correlation is not surprising, but it is not clear to me if the problem was really SSFS or something before SSFS. Experts can better study the data.

The unlocks #1 and #4 come together with a sudden and large fluctuation of the dark fringe (fig 2). A precursor is well visible in BS AA signals (B1p 50MHz QD). The glitches are quite weird: they look like saturations. Before the glitches, the ITF looks quite stable and I cannot find any obvious reason of such a behaviour.

In the unlocks #3 and #5, B1 seems the first to die (fig 3). B1_PD audio channel (fig 4) passes from a normal value to the saturation in 1 ms. Also in this case, the event happens when the ITF looks stable. In fig 5 a normal B1p is visible for 100 s before the unlock. I have checked some other longitudinal and angular error signal, whitout finding anything wrong.

Images attached to this report
Comments to this report:
mwas - 10:37 Wednesday 15 May 2024 (64262) Print this report

Figure 1 and 2. The glitches affecting BS TY are present only on B1p QD1 50MHz and not on B1p QD2 50MHz, so it is not an interferometer problem. They are also not present on B1p QD1 56MHz, so it is not a quadrant hardware problem that should affect all demodulation frequencies. These glitches are very short, ~1ms or less. They affect both H, V and Sum channels.

Figure 3 Looking further they correspond to the phase_noise_corr jumping back and forth by 11 radians. If I am reading well where it comes from in Acl, this is due to the B1s 112MHz phi also jumping by 11 radians. I don't understand why the B1s 112MHz phase jumps.

Images attached to this comment
Search Help